# From Restless to Contextual: A Thresholding Bandit Reformulation for Finite-horizon Improvement

### v4

@@ G.2 Cumulative Reward
  • RRMABB: It uses the explore-then-commit framework. It first randomly gives action and constructs the confidence interval of the unknown transition. After the exploration period, it uses the estimated transition to calculate the OMR policy.
@@
We also note that in the original RRMABB algorithm, it assumes that they can reinitialize at any state for arbitrary number of times. This makes their exploration criteria, which will only stop when sufficient number of samples are collected for each agent, state, action pair, possible. We think this assumption is not impractical and do not use this structure. Therefore, we also change the algorithm so that the exploration period will end after the same number of time-steps instead of number of samples.
Choice of $\gamma$
In this paragraph, we will compare $\epsilon$-GT with different threshold with 2 states. Specifically, we will choose $\gamma\in\{-0.9,0.9\}$. 0.9-0.9 denotes a more moderate choice of $\gamma$ which balances between satisfying Assumption 4.2 and maximizing the reward and $0.9$ denotes a more aggressive choice of $\gamma$ which aims at maximizing the reward. As shown in Figure G.9, $\gamma=-0.9$ outperforms $\gamma=0.9$ across all instances.
Figure G.9: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Two states.
@@
Figure G.10: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Ten states.
In this paragraph, we increase the size of the state space to 10 states and do not impose the stochastic dominance assumption to randomly generate 50 instances to test the performance of $\epsilon$-GT comparing to the benchmarks. Similarly, we also test two settings: 1) a noiseless case where the reward is deterministic, and 2) a noisy case where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 0.1. We would like to note that since UCWhittlePv and RRMABB behave like random policy for 2 states, we do not include that in the 10 states setting.
@@ G.3 Ablation Study
In this section, we will test the performance of Alg. 1 with different choices of $\eta$. Specifically, we will choose four choices of $\eta\in\{0.1,0.5,1,1.5\}$. We will use the same experiment setup as in Section 5 and fix $\gamma=-0.9$.
Figure G.11: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Two states.
As shown in Figure G.11, Alg. 1 is stable over different choices of $\eta$.

### v5

@@ G.2 Cumulative Reward
[…]
@@
[…]
@@
Figure G.10: Average (over time) cumulativere reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward. Ten states.
In this paragraph, we increase the size of the state space to 10 states to randomly generate 50 instances to test the performance of $\epsilon$-GT comparing to the benchmarks. Similarly, we also test two settings: 1) a noiseless case where the reward is deterministic, and 2) a noisy case where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 0.1. We would like to note that since UCWhittlePv and RRMABB behave like random policy for 2 states, we do not include that in the 10 states setting.
@@ G.3 Ablation Study
In this section, we will test the performance of Alg. 1 with different choices of $\eta$ and $\gamma$.
Choice of $\eta$:
We will choose four choices of $\eta\in\{0.1,0.5,1,1.5\}$. We will fix the $\gamma$ and $p$ to align with the choice of $\gamma$ used for experiments in Section 5. As shown in Figures G.11-G.13
Figure G.11: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 binary state instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward.
Figure G.12: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 10 state instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward.
Figure G.13: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 100 repetitions of real dataset. Top: noiseless reward; bottom: noisy reward.
Alg. 1 is stable over different choices of $\eta$ for all of the synthetic binary state, ten state instances and real instance we have tested.
Choice of $\gamma$
In this paragraph, we will compare $\epsilon$-GT with different threshold. We will fix other hyper-parameters to align with choices used in Section 5. As shown in Figure G.14-Figure G.16, for all instances, as $\gamma$ increases, the convergence rate of Alg. 1 is slower. This translates to a lower average reward. Also, we observe that among all instances, a low threshold that satisfies Assump. 4.2 will achieve the best performance.
Figure G.14: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 binary state instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward.
Figure G.15: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 ten state instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward.
Figure G.16: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 100 repetitions of real dataset. Top: noiseless reward; bottom: noisy reward.